# Prospects for PBR detection of KM3-230213A-like events

### v3

@@ 3 Possible neutrino origins of KM3-230213A
To elucidate the origin of KM3-230213A it is pivotal to examine consistency with multi-messenger observations; particularly, constraints from $\gamma$-ray telescopes. It is of common knowledge that transparent astrophysical sources of extreme energy neutrinos would also produce a bright $\gamma$-ray signal. To date, no convincing evidence of $\gamma$-rays accompanying KM3-230213A has been observed [28]. As a consequence, constraints can be placed on a combination of the source redshift and the intergalactic magnetic field strength between the source and Earth [29, 30]. If the strength of the intergalactic magnetic field were $B\gtrsim 3\times 10^{-13}~{\rm G}$, then the $\gamma$-ray emission from farway sources could be attenuated en route to Earth, and the model of transient sources described in [26] would be consistent with observations. On the other hand, the absence of a $\gamma$-ray signal associated to KM3-230213A challenges scenarios in which this neutrino has a Galactic origin, e.g., from the decay of super-heavy dark matter clustered in the halo. Nevertheless, it was pointed out in [31] that if the dark matter were just ordinary matter locked inside primordial black holes (PBHs), then the KM3NeT-ARCA event might have originated from a Galactic transient explosion of a PBH near the end of its evaporation lifetime. Actually, the $\gamma$-ray constraints could be easily accommodated in scenarios with large extra dimensions, because the emission of higher-dimensional PBHs is dominated by Kaluza-Klein modes [32, 33]. It was recently pointed out in [34] that a cosmological population of microscopic black holes, which rapidly evaporate via Hawking radiation, emitting intense, short-lived bursts of neutrinos with $E_{\nu}\gtrsim 10^{8}{\rm GeV}$ can accommodate the groundbreaking KM3NeT-ARCA detection while remaining consistent with IceCube’s non-detections. These microscopic black holes are expected to be produced via bubble collisions generated in the merger of astrophysical black holes; therefore, exploring another window of multi-messenger astrophysics driven by neutrinos and gravitational waves.

### v4

@@ 3 Possible neutrino origins of KM3-230213A
To elucidate the origin of KM3-230213A it is pivotal to examine consistency with multi-messenger observations; particularly, constraints from $\gamma$-ray telescopes. It is of common knowledge that transparent astrophysical sources of extreme energy neutrinos would also produce a bright $\gamma$-ray signal. To date, no convincing evidence of $\gamma$-rays accompanying KM3-230213A has been observed [28]. As a consequence, constraints can be placed on a combination of the source redshift and the intergalactic magnetic field strength between the source and Earth [29, 30]. If the strength of the intergalactic magnetic field were $B\gtrsim 3\times 10^{-13}~{\rm G}$, then the $\gamma$-ray emission from farway sources could be attenuated en route to Earth, and the model of transient sources described in [26] would be consistent with observations. On the other hand, the absence of a $\gamma$-ray signal associated to KM3-230213A challenges scenarios in which this neutrino has a Galactic origin, e.g., from the decay of super-heavy dark matter clustered in the halo. Nevertheless, it was pointed out in [31] that if the dark matter were just ordinary matter locked inside primordial black holes (PBHs), then the KM3NeT-ARCA event might have originated from a Galactic transient explosion of a PBH near the end of its evaporation lifetime. Actually, to accommodate the $\gamma$-ray constraints one should advocate for scenarios with large extra dimensions, because PBHs living in the higher-dimensional bulk would radiate SM singlets that eventually oscillate into active neutrinos [32, 33]. It was recently noted that a cosmological population of microscopic black holes, which rapidly evaporate via Hawking radiation, emitting intense, short-lived bursts of neutrinos with $E_{\nu}\gtrsim 10^{8}{\rm GeV}$ can accommodate the groundbreaking KM3NeT-ARCA detection while remaining consistent with IceCube’s non-detections [34]. These microscopic black holes are expected to be produced via bubble collisions generated in the merger of astrophysical black holes; therefore, exploring another window of multi-messenger astrophysics driven by neutrinos and gravitational waves.